OpenStudy (anonymous):

When I swing at a nail, I drive it all the way in with probability 1/2. With probability 1/4, I hit it half-way in, and with 1/4 probability I miss it entirely. Hitting it half-way in twice will drive it all of the way in. I'm pretty sure that if I swing 4 times at a nail, I'll get it all the way in almost all the time. Let's see if I'm right. How many sequences of 4 swings could leave the nail still not knocked all the way in?

OpenStudy (anonymous):

another question: When I swing at a nail, I drive it all the way in with probability 1/2. With probability 1/4, I hit it half-way in, and with 1/4 probability I miss it entirely. I'm pretty sure that if I swing 4 times at a nail, I'll get it all the way in almost all the time. Let's see if I'm right. What is the probability I fail to get the nail driven completely in?

OpenStudy (welshfella):

first question you could miss the nail 4 times - that's one sequence you might hit it only once in the 4 swings - that's 4 sequences that's a total of 5

OpenStudy (welshfella):

second part probability you miss 4 times = (1/4)^4 P(hitting first time then miss next 3) = (1/4)(1/4)^3 = (1/4)^4 there are 4 different ways of doing this which gives a probability of 4*(1/4)^4 So P(failing to drive the nail in) = 5*(1/4)^4 = 5/256